OpenStudy (anonymous):

Find an equation of a line that goes through (-4,-2) and has a slope =0.

OpenStudy (anonymous):

y=mx + b m=slope b= y-intercept y=0x + b -2= (-4)0 + b b=-2 y=-2

OpenStudy (jlastino):

use the slope point formula y-y1=m(x-x1) y1= -2 x1= -4 m=0